When x number of units are sold, the price of each unit (in
dollars) is given by P=-x/2+75. Find the unit price when the following quantities
are sold: 2, 7, 9, 11.

Answer by checkley71(8403)

P=-X/2+75
P=-2/2+75
P=-1+75=74
--------------
P=-7/2+75
P=-3.5+75
P=71.5
---------------
P=-9/2+75
P=-4.5+75
P=70.5
-----------------------
P=-11/2+75
P=-5.5+75
P=69.5